OpenStudy (gabylovesyou):

Sara selects two cards at random from a standard deck of 52 cards. Which of the following could be used to calculate the probability that she will select two numbered cards if she does not replace the first card before selecting the second? Note: For this problem, face cards and aces are not numbered cards. 36/52 x 35/52 36/52 x 36/51 36/52 x 36/52 36/52 x 35/51

OpenStudy (gabylovesyou):

please help me :D

OpenStudy (anonymous):

The last one: There a nine numbered cards in each suit (2,3,4,5,6,7,8,9,10), and there are four suits, so the total number of numbered cards in a deck is \[9 \times 4 = 36\] So, the probability of the first card being numbered is 36/52, as there are 52 cards in the whole deck. Since you have taken a numbered card out, there are now only 35 numbered cards in the deck, and 51 in total. So the probability of the second card being numbered is 35/51. Therefore, the probability of the two cards being numbered is \[\left(\begin{matrix}36 \\ 52\end{matrix}\right) \times \left(\begin{matrix}35 \\ 51\end{matrix}\right)\]

OpenStudy (gabylovesyou):

Thank you!!

OpenStudy (anonymous):

No problem! :)